OpenStudy (anonymous):

Find and simplify expressions for f(g(x)) and g(f(x)) where f(x)=e^(x)+3 and g(x)=ln(x-3). Also find the inverses. You should get for answers. Help please!

OpenStudy (anonymous):

f(g(x)) = e^(ln(x-3)) + 3 = x - 3 + 3 = x g(f(x)) = ln(e^(x) + 3 - 3) = ln(e^x) = x hence they are inverses of each other

OpenStudy (anonymous):

I'm sorry for the mix up but it's actually: f(x)=e^(x+3) and g(x)=ln(x-3) 1. f(g(x))= 2. g(f(x))= 3. Inverse f(x)= 4. Inverse g(x)= Thank you so much